solve by first eliminating the decimal numbers.
0.5x + 3.4 = 0.2x + 12.4

Answer by khwang(438)

If a^2x-3 * b^2x=a^6-x *b^5x prove that 3 log a =x log a/b
Apply log on both sides we have

(2x-3) log a + 2x log b = (6-x) log a + 5x log b.
(3x - 9) log a = 3x log b.
3x (log a - log b) = 9 log a,
So, log a/b = 3 log a.

Note: as a good habit, you should group terms clearly as:
a^(2x-3) * b^(2x)=a^(6-x) *b^(5x)
kENNY